2) A year ago, you purchased a $1,000 face value bond for $1024. A year later you sold the bond for $1,007 after receiving a coupon payment of $55. What was your rate of capital gain? (Answer in tenth of a percentage i.e. 3.7 & include a minus sign if it is a negative return).
Answer #1

Purchase price = 1024$

Stock price at which is sold = 1007$

Coupon payment = 55$

Total amount after first year = 1007+55 = $1062

%Capital gain = ((1062/1024)-1)*100

% Capital gain = (1.037-1)*100

% Capital gain = 3.7%

The capital gain is 3.7% which is a positive return.